OpenStudy (anonymous):

Adrian has a square bedroom. The area of the room is 169ft^2. What is the length of one side of his bedroom? 6 1/2 ft 13ft 42 1/2 ft 84 1/2 ft will become a fan and will give a medal :)

OpenStudy (anonymous):

@Rushwr

OpenStudy (rushwr):

What do u think ?

OpenStudy (anonymous):

tbh i'm not really sure.

OpenStudy (rushwr):

Since it's a square width = length Area = width *length right? for a square area = width * width \[Area = width ^{2}\] \[width = \sqrt{area} = \sqrt{169} =13ft\]

OpenStudy (rushwr):

width = length so no difference

OpenStudy (rushwr):

any doubts?

OpenStudy (anonymous):

i still don't understand it.

OpenStudy (rushwr):

which part

OpenStudy (rushwr):

area = width * length right?

OpenStudy (rushwr):

in a square width = length So we can write area = width * width OR area = length * length right?

OpenStudy (anonymous):

area = a^2 (a being the side length) so in order to find the side length you would do the square root of 169. The side length would be 13.

OpenStudy (rushwr):

Imagine that we have a square with an unkonw side length "X' Then area = X*X \[Area = X ^{2}\] \[X = \sqrt{Area} = \sqrt{169} = 13\]

OpenStudy (rushwr):

@wormie1999 see if u get it now ?

OpenStudy (anonymous):

i got it. thanks.